OpenStudy (anonymous):

Mitch has $4.55 in dimes and quarters on his desk. He has 7 more quarters than dimes. find the number he has of each coin using consecutive integers.

OpenStudy (lgbasallote):

let x = number of dimes; y = number of quarters 10x + 25y = 455 cents y = 7+x 10x + 25y = 455 (-x + y = 7) 10 10x + 25y = 455 -10x + 10y = 70 35y = 525 y = 15 x = 8 there are 8 dimes and 15 quarters I'm sorry i can only do it using systems of equations